OpenStudy (shubhamsrg):

since many were posting some sort of a tutorial,,i decided to post one(not to be in the competition but because i like the concept :)) this is a very short tutorial which details with a shortcut to mutiply nos. with 99;999;9999;99..... etc

OpenStudy (shubhamsrg):

So here we go (first the method,then the explanantion) : 1)suppose you're multiplying 999 and abc 2)subtract 1 from abc -->these form 1st 3 digits of your product 3)subtract each digit from 9 and write em down from left to right,,which forms 4th,5th and 6th digits 4) add 1 to the result <--you get your ans.. eg. 999 * 748 first 3 digits are (748-1) = 747 next 3 are (9-7),(9-4),(9-8) = 2,5,1 so we have 747251 thus,,the product would be 747251 + 1 = 747252 <--you may verify using calculator next up is the explanation

OpenStudy (shubhamsrg):

the no. is abc now we know that 999 = 1000-1 so ultimately what we are doing is abc * (1000-1) = abc000 - abc so you must be able to see this : a b c 0 0 0 - a b c _____________________ (abc-1)(9-a)(9-b)(10-c) -------------------- and therefore our method works 1st 3 digits are abc-1 next 3 are 9-a,9-b and (9-c +1) the method should take generally a couple of secs to find the product or to get to the ans.. next up are few imp points :

OpenStudy (shubhamsrg):

1)the no. with which you are multiplying 999 with has to be smaller than 999 i.e. 4 digits or greater digit nos. are not part of the trick (ofcorse) 2) same method can be applied for 99,9999,99999... etc ,,but just make sure the no. is less than 99,9999,,etc eg. 9999 * 1453 = (1453 -1);(9-1);(9-4);(9-5);(9-3 +1) = 14528547 3) if you;re multiplying 999 with some 2 digit or 1 digit no..just assume that no. to be 3 digit i.e for e.g 999 * 72 ,,take 72 as 072..so we then have ans as (72-1);(9-0);(9-7);(9-2 +1) = 71928 same with 99,and others.. I'D GUESS THAT'S ALL FOR THIS ONE!! HOPE THAT WAS HELPFUL :)

OpenStudy (anonymous):

yepp very cool...vedic maths..

OpenStudy (shubhamsrg):

^_^

OpenStudy (anonymous):

http://www.youtube.com/watch?v=ZHu0ryjnkYw

OpenStudy (shubhamsrg):

the copied me eh ? lol.. but does he tell you the explanantion?? <ding><ding> !! :P

OpenStudy (shubhamsrg):

i meant he** instead of the**

OpenStudy (shubhamsrg):

:P

OpenStudy (anonymous):

Just use a calculator?

OpenStudy (shubhamsrg):

hmmn,,you may.. but i am pretty sure this'll give quicker results than a calculator,.it helped me atleast

OpenStudy (anonymous):

Just seems like this is unnecessary information to know.

OpenStudy (shubhamsrg):

alright,,i wont show any offence sir..the thought processes are different..hmm

OpenStudy (anonymous):

I just think that knowing formulas for other things and theorems is better than memorizing this. My professor even says that there are some things better worth memorizing than others.

OpenStudy (shubhamsrg):

you find this difficult to memorize? o.O (i agree this will rarely be of any use in any exam,,but still,,its a good trick and i'll stick to it)

OpenStudy (anonymous):

No it's completely useless in real life applications.

OpenStudy (shubhamsrg):

lol..alright,,alright,,you're getting carried away,,calm down sir :D

OpenStudy (anonymous):

Dude you asked if this was hard to memorize when I never get that impression. You basically called me an idiot.

OpenStudy (shubhamsrg):

you basically thought that,,or should i say assumed that sir,,i know you're quite good with the subject..i dont dare call you an idiot buddy..

OpenStudy (shubhamsrg):

you definitely must be right in what you are saying and i have no problems with that :|